OpenStudy (collegekid):

simplify 2+1/2 divided by 3-6/7

OpenStudy (anonymous):

2+1/2=5/2 3-6/7=2+1/7 or 15/7 Multiple the numerator by the reciprocal of the denominator, (5/2)*(7/15)=7/6 Answer: 7/6